find the least positive integral solution of x and y in 396x-763y=12

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your questions below
Given 396x-763y=12
x-79(12)=763t and y-41(12)=396t
for some integer t. This gives all solutions, and the
smallest solution in natural numbers x, y is obtained by
taking t = -1 and turns out to be (185,96).